# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 2. Results
Finally, we have the following technical assumption for the function $\mathcal{K}_{\alpha}$, as in (2.17).
Assumption $\bm{\mathcal{K}_{\alpha}}$. For $\lambda^{*}>0,d^{*}\geq 0$, there exists a function $r\colon(0,\infty)\to\mathbb{R}_{+}$ such that
@@
Discussion, Examples, and Related work. The results presented in this section identify the behaviour of the PAVD model in the ‘infinite lifetime’ and ‘finite lifetime’ with regards to persistence of the maximum degree (or $m^{\mathrm{th}}$-largest degree). In the former regime, we require fairly mild conditions only. In fact, we can prove some of the results in a more general set-up as well, see Section 4. In the ‘finite lifetime’ regime, stronger assumptions are required compared to the ‘infinite lifetime’ regime. This is due to the fact that we need very precise asymptotic expansions of the random variables $I_{n}$ and $O_{n}$, which in this regime no longer converge without rescaling. In particular, we require control over the growth rate of a continuous-time branching process that we analyse to prove the results in Theorem 2.6 (see Section 3 for more details). In Section 6 we discuss some open problems related to this regime.
@@
[…]
@@
Assumptions. Let us then provide some examples that satisfy the assumptions used in the two main theorems. First, we discuss Assumption $\mathcal{K}_{\alpha}$. Here, with $c\in(0,\lambda^{*}/(\lambda*+d^{*}))$, set
$\varepsilon_{t}:=\sup_{i\geq ct}|d(i)-d^{*}|.$ (2.24)
@@
$x_{1}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t\big{)}\qquad\text{and}\qquad x_{i}(t):=\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x_{i-1}(t)\big{)}\qquad\text{for }i>1.$ (2.26)
Then, take $x(t):=x_{k}(t)$ and $r(t):=\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x(t)$. It can be shown, using (the proof of) Lemma 5.5 $(c)$ that this choice of $r$ satisfies Assumption $\mathcal{K}_{\alpha}$. We observe that $\mathcal{K}_{\alpha}(t)=o(t)$ by Lemma 5.5 $(c)$, so that $\varepsilon_{t}=t^{-\beta}$ for some $\beta>0$ satisfies (2.25) for any $k>\beta^{-1}$. That is, as long as $d$ converges to $d^{*}$ at a polynomial rate, Assumption $\mathcal{K}_{\alpha}$ is satisfied. We expect that much weaker conditions on $d$ are sufficient as well, when making more assumptions about the regularity of $b$.
@@
  • $\mathbb{E}\left[D\right]=\infty$, $b$ tends to infinity, $C_{1}:=\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$, and $\widehat{\mu}(C_{1})>1$.
@@
Recent work of Heydenreich and the author in [23] also studies the PAVD, but focusses on lack of persistence, in the sense of (LP). This complements the results presented here and combined these two works provide a clear understanding of the different regimes that can be observed, see Table 2. Further extensions can still be made, however, in particular by not requiring that the death sequence $d$ converges to a constant $d^{*}\geq 0$. We discuss this in more detail in Section 6.
@@ 3. Methodology: Embedding in CMJ branching process
To prove the results presented in Section 2, we make use of a technique where one ‘embeds’ a random discrete process into a continuous-time branching process, a Crump-Mode-Jagers (CMJ) branching process to be exact. The random discrete structure is then equal in distribution to the CMJ branching process, when viewed at certain stopping times, whilst the CMJ branching process in continuous time provides more analytical advantages compared to the discrete process.
General CMJ branching processes. We first introduce some definitions and notation. We let $\mathcal{U}_{\infty}$ denote the Ulam-Harris tree. That is,
$\mathcal{U}_{\infty}:=\{\varnothing\}\cup\bigcup_{k=1}^{\infty}\mathbb{N}^{k}.$ (3.1)
@@
$L^{(v)}:=\sum_{i=1}^{D^{(v)}+1}X(vi).$ (3.2)
Furthermore, we can recursively construct all birth times $\mathcal{B}(v)$ by setting $\mathcal{B}(\varnothing):=0$ and
$\mathcal{B}(vj):=\begin{cases}\mathcal{B}(v)+\sum_{i=1}^{j}X(vi)&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{1,\ldots,D^{(v)}\},\\ \infty&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{D^{(v)}+1,\ldots\}.\end{cases}$ (3.3)
That is, given that $v$ is born in finite time, each child $v1,\ldots,vD^{(v)}$ is born finite time and children $vj$ for $j>D^{(v)}$ are never born. The individual $v$ then dies after $L^{(v)}$ time. If $v$ is never born (i.e. $\mathcal{B}(v)=\infty$) then it is clear that all descendants of $v$ are never born either.
We can then construct the branching process $(\mathrm{BP}(t))_{t\geq 0}$ by setting
$\mathrm{BP}(t):=\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t\}\qquad\text{for }t\geq 0.$ (3.4)
In words, $\mathrm{BP}(t)$ contains all individuals that are born by time $t$. It is clear that individuals $v=v_{1}\cdots v_{m}$ with $v_{1},\ldots,v_{m}\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ are born in the process $\mathrm{BP}$ if and only if $v_{i}\leq D^{(v_{1}\cdots v_{i-1})}$ for all $i\in[m]$. All other individuals are never born. We further define
$\mathcal{A}^{\mathrm{ct}}_{t}:=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\leq t,\sigma_{u}+L^{(u)}>t\},\qquad t\geq 0,$ (3.5)
@@
We can view the point process $\mathcal{R}$ in (2.6), its density $\mu$ in (2.7), and the Laplace transform $\widehat{\mu}$ of $\mu$ in (2.8) from this more general set-up as well. That is, here we set
$\mathcal{R}:=\sum_{k=1}^{D}\delta_{\sum_{i=1}^{k}X(i)}.$ (3.7)
With $\mu(t)\mathrm{d}t=\mathbb{E}\left[\mathcal{R}(t)\mathrm{d}t\right]$ and using Fubini’s theorem, the Laplace transform $\widehat{\mu}$ then equals
$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (3.8)
Persistence in BP. We can view the concepts of persistent $m$-hubs, and persistence and the lack thereof from a continuous-time perspective as well. To this end, we define
$\deg_{t}(u):=\min\bigg{\{}\sup\bigg{\{}k\in\mathbb{N}:\sum_{i=1}^{k}X(ui)\leq t-\mathcal{B}(u)\bigg{\}},D^{(u)}\bigg{\}},\qquad u\in\mathcal{U}_{\infty},t\geq 0,$ (3.9)
as the number of children $u$ has produced by time $t$. We use the convention that we set $\deg_{t}(u)=-\infty$ when $t<\mathcal{B}(u)$, i.e. when $u$ has not (yet) been born. We then define
$O_{t}^{\mathrm{ct}}:=\!\!\min_{v\in\mathcal{A}_{t}^{\mathrm{ct}}}\!\!\mathcal{B}(v)$ (3.10)
as the birth-time of the oldest alive individual at time $t$ as the continuous-time counterpart of $O_{n}$, as in (2.2). Similarly, we define
$I_{t}^{(1),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\big{\}},$ (3.11)
and for $m\geq 2$ fixed,
$I_{t}^{(m),\mathrm{ct}}:=\min\big{\{}\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]}\big{\}}.$ (3.12)
Here $I_{t}^{(m),\mathrm{ct}}$ denotes the birth-time of the oldest alive individual that has the $m^{\mathrm{th}}$ largest number of children at time $t$ and is the continuous-time counterpart of $I_{n}^{(m)}$. For a branching process $(\mathrm{BP}(t))_{t\geq 0}$ we say that it contains a persistent $m$-hub when there exists a random variable $I^{(m),\mathrm{ct}}$ such that
[…]
When $(\mathrm{BP}(t))_{\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, we say that persistence respectively lack of persistence occurs in $(\mathrm{BP}(t))_{t\geq 0}$ when
$(I^{(1),\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t})_{t\geq 0}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$,}$ (3.14)
or
Again, persistence in $(\mathrm{BP}(t))_{t\geq 0}$ is implied when $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $1$-hub, whereas lack of persistence implies that $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$.
The continuous-time preferential attachment with vertex death branching process. We now provide a specific case for the general definition above that we coin the continuous-time preferential attachment with vertex death branching process (CTPAVD), which is directly related to the PAVD model. Recall the random variables $D$, $(E_{i})_{i\in\mathbb{N}_{0}}$, and $(S_{k})_{k\in\mathbb{N}}$ from (2.6). For each individual $v\in\mathcal{U}_{\infty}$, we let $D^{(v)}$ be an i.i.d. copy of $D$, an let each of its inter-birth times $X(vi)$ be an i.i.d. copy of $E_{i-1}$ for each $i\in\mathbb{N}$. It follows that, for each $k\in\mathbb{N}$,
$\sum_{i=1}^{k}X(vi)\overset{\mathrm{d}}{=}S_{k}\qquad\text{and, in particular,}\qquad L^{(v)}\overset{\mathrm{d}}{=}S_{D+1}.$ (3.16)
Continuous-time embedding of PAVD into CTPAVD.
Let us define the following quantities. First, we introduce the characteristics
$\chi_{\mathrm{b}}(t):=\mathbbm{1}_{\{t\geq 0\}},\qquad\text{and}\qquad\chi_{\mathrm{a}}(t):=\mathbbm{1}_{\{0\leq t<L\}},$ (3.17)
@@
$Z_{\square}(t):=\sum_{u\in\mathcal{U}_{\infty}}\chi_{\square}(t-\mathcal{B}(u)).$ (3.18)
@@
$N(t):=Z^{\mathrm{b}}_{t}+(Z^{\mathrm{b}}_{t}-Z^{\mathrm{a}}_{t})\qquad\text{for }t\geq 0,$ (3.19)
as the number of births and deaths that have occurred in the branching process up to time $t$. We define the stopping times $(\tau_{n})_{n\in\mathbb{N}}$ as
$\tau_{n}:=\inf\{t\geq 0:N(t)=n\},\qquad n\in\mathbb{N}.$ (3.20)
@@
The proposition follows from the memoryless property of exponential random variables and properties of minima of exponential random variables. The technique of embedding evolving discrete structures was originally pioneered by Athreya and Karlin [1] and has found fruitful applications in a wide variety of discrete (combinatorial) models, such as Pólya urns, discrete randomly growing trees such as preferential attachment trees and uniform attachment trees, evolving simplicial complexes, and many more.
The description of the branching process $\mathrm{BP}$ in which we embed the PAVD tree is used by Deijfen [13] to determine an explicit description for the limiting offspring distribution of alive individuals. The description of the branching process used here is somewhat different (though equivalent), but provides an analytical advantage. In particular, defining the offspring and lifetime of an individual as an i.i.d. copy of $D$, see (2.6), and $L$, defined in (3.2), respectively, is a perspective which allows us to analyse these quantities to a greater extent.
@@
where $W$ is a non-negative random variable, such that $\mathbb{P}\!\left(W>0\,|\,\mathcal{S}\right)=1$ and hence that $\tau_{n}-\frac{1}{\lambda^{*}}\log n$ converges almost surely. As a result, we can, approximately, relate the quantities in (3.22) via
@@
4. Persistent hubs in general CMJ branching processes
This section is devoted to proving results regarding the (non-)existence of persistent hubs for a general family of CMJ branching processes. In particular, we verify that this family includes the CTPAVD branching process, so that via Proposition 3.1 we obtain Theorem 2.1.
4.1. General results
We first state several assumptions that are required to prove the general results. Recall the definition of the offspring $D^{(v)}$ and inter-birth times $(X(vi))_{i\in\mathbb{N}}$ of an individual $v\in\mathcal{U}_{\infty}$ after (3.1) We view these as i.i.d. copies of random variables $D$ and $(X(i))_{i\in\mathbb{N}}$. Also recall the branching process $(\mathrm{BP}(t))_{t\geq 0}$ in (3.4). We have the following assumption.
Assumption 4.1.
The random variables $D$ and $(X(i))_{i\in\mathbb{N}}$ satisfy the following:
  1. (1)
    $D$ and $(X(i))_{i\in\mathbb{N}}$ are mutually independent.
  2. (2)
    $X(i)<\infty$ for all $i\in\mathbb{N}$ almost surely.
  3. (3)
    $\sum_{k=1}^{\infty}\mathbb{P}\!\left(D\geq k\right)\prod_{i=1}^{k}\mathbb{P}\!\left(X(i)=0\right)<1$.
  4. (4)
    $\mathbb{P}\!\left(D=\infty\right)>0$.
Parts $(1)$ and $(2)$ of Assumption 4.1 are relatively straightforward. Part $(3)$ states that the expected number of children that are born instantaneously (i.e. born at the same time as their parent) is strictly smaller than one, so that in particular sub-trees of the branching process formed by individuals born instantaneously are finite almost surely. Part $(4)$ finally implies that individuals can produce an infinite offspring with positive probability, which is the crucial assumption that distinguishes between Theorems 2.1 and 2.6.
We then have the following result classifying when $(\mathrm{BP}(t))_{t\geq 0}$ does not contains a persistent $m$-hub for any $m\in\mathbb{N}$, and when $(\mathrm{BP}(t))_{t\geq 0}$ does contains a persistent $1$-hub. To this end, we recall the general form of $\widehat{\mu}$ in (3.8).
Theorem 4.2 (Existence of $1$-hubs).
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 is satisfied. We then have:
@@ 4.1. General results
    Suppose the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely. Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub.

### v2

@@ 2. Results
Finally, we have the following technical assumption for the function $\mathcal{K}_{\alpha}$, as in (2.17). Assumption $\bm{\mathcal{K}_{\alpha}}$. For $\lambda^{*}>0,d^{*}\geq 0$, there exists a function $r\colon(0,\infty)\to\mathbb{R}_{+}$ such that
@@
Discussion, Examples, and Related work. The results presented in this section identify the behaviour of the PAVD model in the ‘infinite lifetime’ and ‘finite lifetime’ with regards to persistence of the maximum degree (or $m^{\mathrm{th}}$-largest degree, in general). In the former regime, we require fairly mild conditions only. In the ‘finite lifetime’ regime, stronger assumptions are required compared to the ‘infinite lifetime’ regime. This is due to the fact that we need very precise asymptotic expansions of the random variables $I_{n}^{(1)}$ and $O_{n}$, which in this regime no longer converge without rescaling. In particular, we require control over the growth rate of a continuous-time branching process that we analyse to prove the results in Theorem 2.6 (see Section 3 for more details). In Section 6 we discuss some open problems related to this regime.
@@
$({}$ $\text{RaO})$ $b$ $=(1,2,3,\ldots)$ $\text{and }d=(1,2,\tfrac{3}{2},\tfrac{3}{2},\ldots),$ (2.23)
$({}$ $\text{RdY}1)$ $b$ $=(1,2,3,\ldots)$ $\text{and }d=(\tfrac{1}{4},2,\tfrac{3}{2},\tfrac{3}{2},\ldots),$
$({}$ $\text{RdY}2)$ $b$ $=(\tfrac{1}{4},2,3,\ldots)$ $\text{and }d=(1,2,\tfrac{3}{2},\tfrac{3}{2},\ldots).$
@@
Assumptions. Let us then provide some examples that satisfy the assumptions used in the two main theorems. First, we discuss Assumption $\mathcal{K}_{\alpha}$. Here, with $c\in(0,\lambda^{*}/(\lambda^{*}+d^{*}))$, set
$\varepsilon_{t}\coloneq\sup_{i\geq ct}|d(i)-d^{*}|.$ (2.24)
@@
$x_{1}(t)\coloneq\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t\big)\qquad\text{and}\qquad x_{i}(t)\coloneq\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x_{i-1}(t)\big)\qquad\text{for }i>1.$ (2.26)
Then, take $x(t)\coloneq x_{k}(t)$ and $r(t)\coloneq\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-x(t)$. It can be shown, using (the proof of) Lemma 5.5(c)(c) that this choice of $r$ satisfies Assumption $\mathcal{K}_{\alpha}$. We observe that $\mathcal{K}_{\alpha}(t)=o(t)$ by Lemma 5.5(c)(c), so that $\varepsilon_{t}=t^{-\beta}$ for some $\beta>0$ satisfies (2.25) for any $k>\beta^{-1}$. That is, as long as $d$ converges to $d^{*}$ at a polynomial rate, Assumption $\mathcal{K}_{\alpha}$ is satisfied. We expect that much weaker conditions on $d$ are sufficient as well, when making more assumptions about the regularity of $b$.
@@
  • $\mathbb{E}\left[D\right]=\infty$, $b$ tends to infinity, $C_{1}\coloneq\limsup_{i\to\infty}\frac{b(i)}{i}<\infty$, and $\widehat{\mu}(C_{1})>1$.
@@
Recent work of Heydenreich and the author in [23] also studies the PAVD model, but focusses on lack of persistence, in the sense of (LP). This complements the results presented here and combined these two works provide a clear understanding of the different regimes that can be observed, see Table 2. Further extensions can still be made, however, in particular by not requiring that the death sequence $d$ converges to a constant $d^{*}\geq 0$. We discuss this in more detail in Section 6.
@@ 3. Methodology: Embedding in CMJ branching process
To prove the results presented in Section 2, we make use of a technique where one ‘embeds’ a random discrete process into a Crump–Mode-Jagers (CMJ) continuous-time branching process. The random discrete structure is then equal in distribution to the CMJ branching process, when viewed at certain stopping times, whilst the CMJ branching process in continuous time provides more analytical advantages compared to the discrete process.
General CMJ branching processes. We first introduce some definitions and notation to facilitate the introduction of the CMJ branching process that relates to the discrete PAVD model. We let $\mathcal{U}_{\infty}$ denote the Ulam-Harris tree. That is,
$\mathcal{U}_{\infty}\coloneq\{\varnothing\}\cup\bigcup_{k=1}^{\infty}\mathbb{N}^{k}.$ (3.1)
@@
$L^{(v)}\coloneq\sum_{i=1}^{D^{(v)}+1}X(vi).$ (3.2)
Furthermore, we can recursively construct all birth times $\mathcal{B}(v)$ by setting $\mathcal{B}(\varnothing)\coloneq 0$ and
$\mathcal{B}(vj)\coloneq\begin{cases}\mathcal{B}(v)+\sum_{i=1}^{j}X(vi)&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{1,\ldots,D^{(v)}\},\\ \infty&\mbox{for }v\in\mathcal{U}_{\infty},j\in\{D^{(v)}+1,\ldots\}.\end{cases}$ (3.3)
Given that $v$ is born in finite time, each child $v1,\ldots,vD^{(v)}$ is born finite time and children $vj$ for $j>D^{(v)}$ are never born (as their birth-time is infinite). The individual $v$ then dies after $L^{(v)}$ time. If $v$ is never born (i.e. $\mathcal{B}(v)=\infty$) then it is clear that all descendants of $v$ are never born either.
We can then construct a branching process $(\mathrm{BP}(t))_{t\geq 0}$ by setting
$\mathrm{BP}(t)\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t\}\qquad\text{for }t\geq 0.$ (3.4)
It is clear that individuals $v=v_{1}\cdots v_{m}$ with $v_{1},\ldots,v_{m}\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ are born in the process $\mathrm{BP}$ if and only if $v_{i}\leq D^{(v_{1}\cdots v_{i-1})}$ for all $i\in[m]$. All other individuals are never born. We further define
$\mathcal{A}^{\mathrm{ct}}_{t}\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\leq t<\mathcal{B}(u)+L^{(u)}\},\qquad\text{ for }t\geq 0,$ (3.5)
@@
Continuous-time embedding of PAVD. We now provide a specific case for the general definition above that we coin the continuous-time preferential attachment with vertex death branching process (CTPAVD), which is directly related to the PAVD model. Recall the random variables $D$, $(E_{i})_{i\in\mathbb{N}_{0}}$, and $(S_{k})_{k\in\mathbb{N}}$ from (2.6). For each individual $v\in\mathcal{U}_{\infty}$, we let $D^{(v)}$ be an i.i.d. copy of $D$, and set $X(vi)=E_{i-1}^{(v)}$, where $E^{(v)}_{i-1}$ is an i.i.d. copy of $E_{i-1}$ for each $i\in\mathbb{N}$. With $S_{k}^{(v)}\coloneq\sum_{i=0}^{k-1}E^{(v)}_{i}$, we can write for each $v\in\mathcal{U}_{\infty}$ and $k\in\mathbb{N}$,
$\sum_{i=1}^{k}X(vi)=S_{k}^{(v)}\qquad\text{and, in particular,}\qquad L^{(v)}\overset{\mathrm{d}}{=}S_{D+1}^{(v)}.$ (3.7)
Let us introduce for $v\in\mathcal{U}_{\infty}$ the characteristics
$\chi^{(v)}_{\mathrm{b}}(t)\coloneq\mathbbm{1}_{\{t\geq 0\}},\qquad\text{and}\qquad\chi^{(v)}_{\mathrm{a}}(t)\coloneq\mathbbm{1}_{\{0\leq t<L^{(v)}\}},$ (3.8)
@@
$Z_{\square}(t)\coloneq\sum_{v\in\mathcal{U}_{\infty}}\chi^{(v)}_{\square}(t-\mathcal{B}(v)).$ (3.9)
@@
$N(t)\coloneq Z^{\mathrm{b}}_{t}+(Z^{\mathrm{b}}_{t}-Z^{\mathrm{a}}_{t})\qquad\text{for }t\geq 0,$ (3.10)
as the number of births and deaths that have occurred in the branching process up to time $t$. We let $(\tau_{n})_{n\in\mathbb{N}}$ be stopping times, defined as
$\tau_{n}\coloneq\inf\{t\geq 0:N(t)=n\},\qquad n\in\mathbb{N}.$ (3.11)
@@
The proposition follows from the memoryless property of exponential random variables and properties of minima of exponential random variables. The technique of embedding evolving discrete structures was originally pioneered by Athreya and Karlin [1] and has found fruitful applications in a wide variety of discrete (combinatorial) models, such as Pólya urns, discrete randomly growing trees such as preferential attachment trees and uniform attachment trees (as first observed by Pittel [31]), evolving simplicial complexes, and many more.
The description of the branching process $\mathrm{BP}$ in which we embed the PAVD tree is used by Deijfen [13] to determine an explicit description for the limiting offspring distribution of alive individuals. The description of the branching process used here is somewhat different (though equivalent). Namely, the explicit construction in terms of the sum of inter-birth times $S_{k}$, the offspring random variable $D$, and the lifetime $L$, was not used by Deijfen. This viewpoint provides an analytical advantage. In particular, it allows us to analyse the lifetime distribution of individuals and the number of children an individual produces over time to a greater extent.
Persistence in the CTPAVD branching process.
We can view the concepts of persistent $m$-hubs, and persistence and the lack thereof from a continuous-time perspective as well. To this end, we define
$\deg_{t}(u)\coloneq\min\big\{\sup\big\{k\in\mathbb{N}_{0}:S_{k}^{(u)}\leq t-\mathcal{B}(u)\big\},D^{(u)}\big\},\qquad u\in\mathcal{U}_{\infty},t\geq 0,$ (3.13)
as the number of children $u$ has produced by time $t$. We use the convention that $S_{0}^{(v)}=0$ for all $v$ and that we set $\deg_{t}(u)=$-$\infty$ when $t<\mathcal{B}(u)$, i.e. when $u$ has not (yet) been born. We also define
$O_{t}^{\mathrm{ct}}\coloneq\!\!\min_{v\in\mathcal{A}_{t}^{\mathrm{ct}}}\!\mathcal{B}(v)$ (3.14)
as the birth-time of the oldest alive individual at time $t$, which is the continuous-time counterpart of $O_{n}$, as in (2.2). Similarly, we define
$I_{t}^{(1),\mathrm{ct}}\coloneq\min\big\{\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\big\},$ (3.15)
and for $m\geq 2$ fixed,
$I_{t}^{(m),\mathrm{ct}}\coloneq\min\big\{\mathcal{B}(v):v\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]},\deg_{t}(v)\geq\deg_{t}(u)\text{ for all }u\in\mathcal{A}^{\mathrm{ct}}_{t}\setminus\{I_{t}^{(i),\mathrm{ct}}\}_{i\in[m-1]}\big\}.$ (3.16)
Here $I_{t}^{(m),\mathrm{ct}}$ denotes the birth-time of the oldest alive individual that has the $m^{\mathrm{th}}$ largest number of children at time $t$ and is the continuous-time counterpart of $I_{n}^{(m)}$. We say a branching process $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent elder when there exists a random variable $O^{\mathrm{ct}}$ that is finite almost surely, such that
$O^{\mathrm{ct}}_{t}\to O^{\mathrm{ct}}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely as $t\to\infty$}.$ (3.17)
We say that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when there exists an almost surely finite random variable $I^{(m),\mathrm{ct}}$ such that
$I_{t}^{(m),\mathrm{ct}}\to I^{(m),\mathrm{ct}}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely as $t\to\infty$}.$ (3.18)
When $(\mathrm{BP}(t))_{\geq 0}$ does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, we say that persistence occurs in $(\mathrm{BP}(t))_{t\geq 0}$ when
[…]
and that lack of persistence occurs when
As in the discrete setting, $(\mathrm{BP}(t))_{t\geq 0}$ containing a persistent $1$-hub implies that it contains a persistent elder, and it implies persistence in $(\mathrm{BP}(t))_{t\geq 0}$. On the other, hand, lack of persistence implies that $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $1$-hub.
@@
where $W$ is a non-negative random variable, such that $\mathbb{P}\!\left(W>0\,|\,\mathcal{S}\right)=1$ and hence that $\tau_{n}-\frac{1}{\lambda^{*}}\log n$ converges almost surely (see also Proposition 5.3). As a result, we can, approximately, relate the quantities in (3.21) via
@@
In Sections 4 and 5 we analyse the CMJ branching process defined in this section and, in particular, prove results under which conditions the behaviour in (3.18) and (3.19) occurs. In each of these sections, we use these results when proving the main theorems by leveraging Proposition 3.1.
4. (Non)-existence of persistent hubs
This section is devoted to proving results regarding the (non-)existence of persistent hubs for the CTPAVD branching process introduced in Section 3. Via Proposition 3.1 we then obtain Theorem 2.1.
We first state the following result classifying when $(\mathrm{BP}(t))_{t\geq 0}$ does (not) contain a persistent $m$-hub for any $m\in\mathbb{N}$
Theorem 4.1 ((Non-)existence of $m$-hubs).
Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7). Assume that Assumption (N-E) is satisfied and Assumption (F-D) is not satisfied.
@@
    If Assumption (C-V) is not satisfied, then for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub.